# GRAPHITE: Graph-Based Interpretable Tissue Examination for Enhanced Explainability in Breast Cancer Histopathology

### v1

@@ 1 Introduction
    Comprehensive Evaluation of Visualisation Performance: GRAPHITE is rigorously evaluated using both standard performance metrics and the newly proposed metrics (ThS and ThR), demonstrating superior results compared to existing attention and gradient based methods across all evaluation criteria.
@@ 3.3.1 Stage 1: MIL-Based Classification
Patch Extraction: TMAs are typically composed of circular cores, each representing a distinct region of interest. At 40$\times$ magnification, a spatial resolution of 0.25 µm per pixel is achieved, which provides sufficient detail for histological analysis of cell structures, nuclei and other tissue features. Each core is systematically divided into non-overlapping patches, with each patch being $224\times 224$ pixels in size, denoted as $\mathbf{x}_{i}\in\mathbb{R}^{224\times 224\times 3}$, where the third dimension represents the RGB channels from the H&E staining. For a given TMA core $k$, the set of patches extracted is represented as $\mathcal{P}_{k}=\{\mathbf{x}_{i}\}_{i=1}^{N_{k}}$, where $N_{k}$ is the total number of patches extracted from core $k$. Each patch is treated as an independent data point, processed through the feature extraction pipeline.
@@
Feature Extraction With ResNet18: ResNet18, a deep CNN, is used as the feature extractor in the first stage of the pipeline [61, 62]. The model we use is pretrained on the NCT-CRC-HE-100K dataset, which includes 100,000 patches of colorectal cancer and normal tissue. The pretrained weights [^1] https://huggingface.co/1aurent/resnet18.tiatoolbox-kather100k provide a strong initialisation, allowing the model to capture general histological features.
@@
Patch-Level Aggregation Using Attention Mechanism: After features are extracted from each patch, they are first transformed into an embedding space using a patch projector. The patch projector consists of a series of fully connected layers (512 and 128 units) that project the high-dimensional patch features into a more compact and discriminative embedding space. This embedding serves as the foundation for further processing in both classification and subsequent visualisation stages.
@@
Core-Level Classification and Patient Projector: Following the attention-based aggregation of patch-level features, the core-level feature vector $\mathbf{z}_{k}$ representing the TMA core is passed through the patient projector, which consists of two fully connected layers (respectively 512 and 128 units). These layers further transform the core-level feature into a more compact and discriminative representation. The transformed feature vector $\mathbf{p}_{j}=f_{\text{dense}}(\mathbf{z}_{k})$ is then passed to the final classifier, which performs binary classification by computing the probability $\hat{y}_{k}\in[0,1]$ that the core $k$ contains tumour tissue. This probability is obtained by applying a sigmoid function $\zeta(\mathbf{W}^{\top}\mathbf{z}_{k}+b)$, where $\mathbf{W}$ and $b$ are the learned weights and bias of the dense layer. The model is trained by minimising the binary cross-entropy loss:
@@ 3.3.2 Stage 2: GRAPHITE-Based Visualisation
Multiscale Patch Extraction: GRAPHITE begins with the extraction of patches from TMA cores at multiple magnification levels, capturing tissue features at varying spatial resolutions. Three magnification levels are employed: Level 0 (40$\times$) at a resolution of 0.25 µm/pixel, Level 1 (20$\times$) at a resolution of 0.50 µm/pixel, and Level 2 (10$\times$) at a resolution of 1.00 µm/pixel. For each magnification level, patches of size $224\times 224$ pixels are extracted. These patches represent specific regions of interest within the TMA core. Higher magnifications capture detailed cellular-level features, while lower magnifications provide a broader contextual view of the surrounding tissue architecture. Let $\mathcal{P}^{m}_{k}=\{\mathbf{x}^{m}_{i}\}_{i=1}^{N^{m}_{k}}$ denote the set of patches extracted from core $k$ at magnification level $m$, where $N^{m}_{k}$ is the number of patches at magnification $m$.
Feature Extraction Using Trained Model: Patch-level features are extracted using a combination of ResNet18 and a patch projector. Each patch is first processed by ResNet18, which outputs a feature vector representing the patch in a high-dimensional feature space. Both the ResNet18 model and the patch projector are fine-tuned during the classification stage (Stage 1) to extract domain-specific histopathology features effectively. The patch-level features are then further refined using the patch projector, which transforms the ResNet18 output into a compact, discriminative embedding (128 dimensions). This embedding facilitates better integration of patch-level features across multiple magnification levels, enhancing the model’s ability to perform multiscale aggregation and analysis. These combined features from ResNet18 and the patch projector serve as the foundation for the subsequent stages of the GRAPHITE framework for improved visualisation.
Hierarchical Graph Construction: The key innovation in GRAPHITE is its use of an hierarchical graph structure to model the spatial and scalewise relationships between patches of the same core at different magnification levels. Each node in the graph corresponds to a patch, and edges between nodes represent spatial relationships, both within and across magnification levels. Formally, let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ represent the hierarchical graph, where $\mathcal{V}$ is the set of nodes and $\mathcal{E}$ is the set of edges. $\mathcal{V}$ is composed of patches $\mathcal{P}_{k}^{m}$ extracted from the TMA core $k$ at magnification level $m$. Thus, the overall set of nodes is defined as:
@@
Graph Attention Network (GAT): The GAT in GRAPHITE plays a crucial role in refining patch-level features by leveraging spatial and cross-scale relationships captured within the hierarchical graph. After constructing the hierarchical graph from patches extracted across multiple magnification levels, GAT aggregates information from neighbouring patches, enabling the model to capture both fine-grained spatial detail and broader contextual dependencies across scales. GAT is composed of two parallel attention mechanisms that process different types of edges: spatial edges that connect patches within the same magnification level, and cross-scale edges that link patches across different magnification levels (Figure 2). Each mechanism employs multihead attention, allowing the model to assign distinct weights to neighbouring nodes and focus on the most relevant spatial and hierarchical relationships.
@@
Scalewise Attention Network (SAN): Following the GAT-based feature enhancement, SAN is employed to learn importance weights specific to each magnification level, enabling the model to integrate multiscale information effectively. For each magnification level $m$ in the hierarchy, SAN computes a level-specific attention score for patches at that level. The network consists of a set of attention modules, each dedicated to a particular magnification level, which outputs attention scores based on the node features.
@@
Saliency Map Generation Using Multiscale and Confidence-Based Fusion: In GRAPHITE, the saliency map generation employs the fusion of attention maps using multiple strategies, including multilevel fusion and confidence-based adaptive fusion. This results in a final saliency map that integrates relevant information from different attention mechanisms (Figure 3).
@@ 3.4.2 Stage 2: Self-Supervised Training With InfoMax and Scalewise Loss
InfoMax Loss: This loss is designed to maximise the mutual information between patch-level features and the core-level (or graph-level) representations. For each magnification level, this loss computes the similarity between the normalised embeddings of patches and the aggregated core-level embedding, promoting feature consistency across different scales. The InfoMax loss is calculated as:
@@
Scalewise Loss: This loss focuses on maintaining consistency and relevance across different magnification levels. It consists of two main components: interscale consistency and intrascale discrimination. Interscale consistency enforces similarity between embeddings from adjacent scales, calculated using the dot product between normalised embeddings of nodes from different scales. Intrascale discrimination aims to differentiate embeddings within the same scale by discouraging similarity among different nodes at the same level. For node embeddings $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ at levels $m$ and $l$, the scalewise loss is defined as:
@@
Total Loss: The total loss for self-supervised training in Stage 2 is defined as the sum of two losses:
@@ 3.5.7 Balanced Accuracy (BA)
Table 1: Performance comparison of different XAI methods.
Method mAP AUROC AUPRC mIoU ThS ThR BA CXPS
GradCAM 0.44 0.86 0.55 0.24 0.17 0.20 0.60 0.48
GradCAM++ 0.50 0.89 0.62 0.31 0.31 0.40 0.63 0.55
AblationCAM 0.44 0.77 0.46 0.25 0.20 0.10 0.59 0.45
EigenCAM 0.45 0.80 0.54 0.20 0.11 0.10 0.58 0.44
FullGrad 0.52 0.91 0.65 0.34 0.36 0.50 0.64 0.58
Attention 0.55 0.92 0.77 0.38 0.43 0.60 0.67 0.62
GRAPHITE-Base 0.55 0.93 0.77 0.37 0.40 0.60 0.67 0.62
GRAPHITE-V1 0.55 0.93 0.76 0.39 0.47 0.60 0.68 0.63
GRAPHITE-V2 0.56 0.94 0.78 0.41 0.50 0.70 0.68 0.65
@@ 3.5.8 Comprehensive XAI Performance Score (CXPS)
Figure 4: Receiver operating characteristic (ROC) curves for various XAI methods on breast cancer TMA analysis. AUROC for each method is displayed in the legend, with GRAPHITE-V2 achieving the highest AUROC of 0.94, followed closely by GRAPHITE-V1 and GRAPHITE-Base, both at 0.93. This performance demonstrates the superior discriminative ability of GRAPHITE variants compared to other XAI methods, such as Attention (AUROC = 0.92) and FullGrad (AUROC = 0.91). The colour gradient along the curves represents different threshold values.
Figure 5: Precision-recall (PR) curves for various XAI methods on breast cancer TMA analysis. AUPRC for each method is displayed in the legend, with GRAPHITE-V2 achieving the highest AUPRC of 0.78, followed by GRAPHITE-Base and Attention, both with an AUPRC of 0.77. This performance highlights the superior PR balance of GRAPHITE variants compared to other XAI methods, such as FullGrad (AUPRC = 0.65) and Grad-CAM++ (AUPRC = 0.62). The colour gradient along the curves represents different threshold values.
@@ 3.7 Comparison of GRAPHITE Variants
Figure 6: Comparison of XAI methods based on the CXPS metric. The GRAPHITE variants (GRAPHITE-Base, GRAPHITE-V1, GRAPHITE-V2) achieve the highest CXPS scores, with GRAPHITE-V2 scoring 0.65, indicating its superior balance of interpretability and predictive performance. Other methods, such as Attention and FullGrad, perform moderately well, while EigenCAM, AblationCAM and GradCAM achieve comparatively lower CXPS scores, reflecting limitations in either interpretability or predictive capability. This evaluation highlights GRAPHITE’s advantage in delivering a reliable and interpretable solution for breast cancer TMA analysis.
@@ 3.8 Comparison With Traditional XAI Methods
Gradient-based methods, such as GradCAM and its advanced variant, Grad-CAM++, which utilize class-specific gradient information, achieved moderate AUROCs of 0.87 and 0.89 respectively, demonstrating their foundational class-discriminative capabilities. FullGrad, which extends gradient-based approaches by incorporating both input and neuron sensitivity components, showed improved performance (AUROC: 0.91) but exhibited limitations in localization precision metrics (mAP: 0.52, mIoU: 0.34). EigenCAM’s principal component analysis approach and AblationCAM’s feature removal strategy demonstrated comparatively lower performance across all metrics, suggesting potential limitations in their underlying theoretical frameworks.
Attention-based methods, like MIL-Attention establishes a competitive baseline (AUROC: 0.92, AUPRC: 0.77, ThR: 0.60), matching GRAPHITE-Base’s performance. GRAPHITE-V1 advances the discriminative capability with an AUROC of 0.93, though with a marginally lower AUPRC (0.76). GRAPHITE-V2 ultimately achieves superior performance across all metrics (AUROC: 0.94, AUPRC: 0.78, ThS: 0.50, ThR: 0.70), highlighting the effectiveness of its integrated architectural innovations in delivering both reliable predictions and interpretable visualizations (Figure 8).

### v2

@@ 1 Introduction
    Comprehensive Evaluation of Visualisation Performance:
    GRAPHITE is rigorously evaluated using both standard performance metrics and the newly proposed metrics (ThS and ThR), demonstrating superior results compared to existing attention and gradient based methods across all evaluation criteria.
@@ 2 Literature Review
Beyond gradient-based methods, model-agnostic techniques like LIME (Local Interpretable Model-agnostic Explanations) [55] and SHAP (SHapley Additive exPlanations) [56] offer alternative approaches to interpretability. LIME explains individual predictions by learning a simpler, interpretable model locally around the prediction instance. SHAP uses a game-theoretic approach to assign contribution values to each feature for a prediction. These methods have been applied in medical imaging, including breast cancer histopathology [57], offering flexibility as they can be applied to any underlying classifier.
@@ 3.3.1 Stage 1: MIL-Based Classification
Patch Extraction
TMAs are typically composed of circular cores, each representing a distinct region of interest. At 40$\times$ magnification, a spatial resolution of 0.25 µm per pixel is achieved, which provides sufficient detail for histological analysis of cell structures, nuclei and other tissue features. Each core is systematically divided into non-overlapping patches, with each patch being $224\times 224$ pixels in size, denoted as $\mathbf{x}_{i}\in\mathbb{R}^{224\times 224\times 3}$, where the third dimension represents the RGB channels from the H&E staining. For a given TMA core $k$, the set of patches extracted is represented as $\mathcal{P}_{k}=\{\mathbf{x}_{i}\}_{i=1}^{N_{k}}$, where $N_{k}$ is the total number of patches extracted from core $k$. Each patch is treated as an independent data point, processed through the feature extraction pipeline.
@@
Feature Extraction With ResNet18
ResNet18, a deep CNN, is used as the feature extractor in the first stage of the pipeline [64, 65]. The model we use is pretrained on the NCT-CRC-HE-100K dataset, which includes 100,000 patches of colorectal cancer and normal tissue. The pretrained weights [^1] https://huggingface.co/1aurent/resnet18.tiatoolbox-kather100k provide a strong initialisation, allowing the model to capture general histological features.
@@
Patch-Level Aggregation Using Attention Mechanism
After features are extracted from each patch, they are first transformed into an embedding space using a patch projector. The patch projector consists of a series of fully connected layers (512 and 128 units) that project the high-dimensional patch features into a more compact and discriminative embedding space. This embedding serves as the foundation for further processing in both classification and subsequent visualisation stages.
@@
Core-Level Classification and Patient Projector
Following the attention-based aggregation of patch-level features, the core-level feature vector $\mathbf{z}_{k}$ representing the TMA core is passed through the patient projector, which consists of two fully connected layers (respectively 512 and 128 units). These layers further transform the core-level feature into a more compact and discriminative representation. The transformed feature vector $\mathbf{p}_{j}=f_{\text{dense}}(\mathbf{z}_{k})$ is then passed to the final classifier, which performs binary classification by computing the probability $\hat{y}_{k}\in[0,1]$ that the core $k$ contains tumour tissue. This probability is obtained by applying a sigmoid function $\zeta(\mathbf{W}^{\top}\mathbf{z}_{k}+b)$, where $\mathbf{W}$ and $b$ are the learned weights and bias of the dense layer. The model is trained by minimising the binary cross-entropy loss:
@@ 3.3.2 Stage 2: GRAPHITE-Based Visualisation
Multiscale Patch Extraction
GRAPHITE begins with the extraction of patches from TMA cores at multiple magnification levels, capturing tissue features at varying spatial resolutions. Three magnification levels are employed: Level 0 (40$\times$) at a resolution of 0.25 µm/pixel, Level 1 (20$\times$) at a resolution of 0.50 µm/pixel, and Level 2 (10$\times$) at a resolution of 1.00 µm/pixel. For each magnification level, patches of size $224\times 224$ pixels are extracted. These patches represent specific regions of interest within the TMA core. Higher magnifications capture detailed cellular-level features, while lower magnifications provide a broader contextual view of the surrounding tissue architecture. Let $\mathcal{P}^{m}_{k}=\{\mathbf{x}^{m}_{i}\}_{i=1}^{N^{m}_{k}}$ denote the set of patches extracted from core $k$ at magnification level $m$, where $N^{m}_{k}$ is the number of patches at magnification $m$.
Feature Extraction Using Trained Model
Patch-level features are extracted using a combination of ResNet18 and a patch projector. Each patch is first processed by ResNet18, which outputs a feature vector representing the patch in a high-dimensional feature space. Both the ResNet18 model and the patch projector are fine-tuned during the classification stage (Stage 1) to extract domain-specific histopathology features effectively. The patch-level features are then further refined using the patch projector, which transforms the ResNet18 output into a compact, discriminative embedding (128 dimensions). This embedding facilitates better integration of patch-level features across multiple magnification levels, enhancing the model’s ability to perform multiscale aggregation and analysis. These combined features from ResNet18 and the patch projector serve as the foundation for the subsequent stages of the GRAPHITE framework for improved visualisation.
Hierarchical Graph Construction
The key innovation in GRAPHITE is its use of an hierarchical graph structure to model the spatial and scalewise relationships between patches of the same core at different magnification levels. Each node in the graph corresponds to a patch, and edges between nodes represent spatial relationships, both within and across magnification levels. Formally, let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ represent the hierarchical graph, where $\mathcal{V}$ is the set of nodes and $\mathcal{E}$ is the set of edges. $\mathcal{V}$ is composed of patches $\mathcal{P}_{k}^{m}$ extracted from the TMA core $k$ at magnification level $m$. Thus, the overall set of nodes is defined as:
@@
Graph Attention Network (GAT)
The GAT in GRAPHITE plays a crucial role in refining patch-level features by leveraging spatial and cross-scale relationships captured within the hierarchical graph. After constructing the hierarchical graph from patches extracted across multiple magnification levels, GAT aggregates information from neighbouring patches, enabling the model to capture both fine-grained spatial detail and broader contextual dependencies across scales. GAT is composed of two parallel attention mechanisms that process different types of edges: spatial edges that connect patches within the same magnification level, and cross-scale edges that link patches across different magnification levels (Figure 2). Each mechanism employs multihead attention, allowing the model to assign distinct weights to neighbouring nodes and focus on the most relevant spatial and hierarchical relationships.
@@
Scalewise Attention Network (SAN)
Following the GAT-based feature enhancement, SAN is employed to learn importance weights specific to each magnification level, enabling the model to integrate multiscale information effectively. For each magnification level $m$ in the hierarchy, SAN computes a level-specific attention score for patches at that level. The network consists of a set of attention modules, each dedicated to a particular magnification level, which outputs attention scores based on the node features.
@@
Saliency Map Generation Using Multiscale and Confidence-Based Fusion
In GRAPHITE, the saliency map generation employs the fusion of attention maps using multiple strategies, including multilevel fusion and confidence-based adaptive fusion. This results in a final saliency map that integrates relevant information from different attention mechanisms (Figure 3).
@@ 3.4.2 Stage 2: Self-Supervised Training With InfoMax and Scalewise Loss
InfoMax Loss
This loss is designed to maximise the mutual information between patch-level features and the core-level (or graph-level) representations. For each magnification level, this loss computes the similarity between the normalised embeddings of patches and the aggregated core-level embedding, promoting feature consistency across different scales. The InfoMax loss is calculated as:
@@
Scalewise Loss
This loss focuses on maintaining consistency and relevance across different magnification levels. It consists of two main components: interscale consistency and intrascale discrimination. Interscale consistency enforces similarity between embeddings from adjacent scales, calculated using the dot product between normalised embeddings of nodes from different scales. Intrascale discrimination aims to differentiate embeddings within the same scale by discouraging similarity among different nodes at the same level. For node embeddings $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ at levels $m$ and $l$, the scalewise loss is defined as:
@@
Total Loss
The total loss for self-supervised training in Stage 2 is defined as the sum of two losses:
@@ 3.5.3 Mean Average Precision (mAP)
Table 1: Performance comparison of different XAI methods.
Method mAP AUROC AUPRC mIoU ThS ThR BA CXPS
GradCAM 0.44 0.86 0.55 0.24 0.17 0.20 0.60 0.48
GradCAM++ 0.50 0.89 0.62 0.31 0.31 0.40 0.63 0.55
AblationCAM 0.44 0.77 0.46 0.25 0.20 0.10 0.59 0.45
EigenCAM 0.45 0.80 0.54 0.20 0.11 0.10 0.58 0.44
FullGrad 0.52 0.91 0.65 0.34 0.36 0.50 0.64 0.58
Attention 0.55 0.92 0.77 0.38 0.43 0.60 0.67 0.62
SHAP 0.36 0.61 0.31 0.15 0.07 0.10 0.52 0.35
LIME 0.45 0.83 0.52 0.25 0.15 0.30 0.60 0.48
GRAPHITE-Base 0.55 0.93 0.77 0.37 0.40 0.60 0.67 0.62
GRAPHITE-V1 0.55 0.93 0.76 0.39 0.47 0.60 0.68 0.63
GRAPHITE-V2 0.56 0.94 0.78 0.41 0.50 0.70 0.68 0.65
@@ 3.5.8 Comprehensive XAI Performance Score (CXPS)
Figure 4: Receiver operating characteristic (ROC) curves for various XAI methods on breast cancer TMA analysis. AUROC for each method is displayed in the legend, with GRAPHITE-V2 achieving the highest AUROC of 0.94, followed closely by GRAPHITE-V1 and GRAPHITE-Base, both at 0.93. This performance demonstrates the superior discriminative ability of GRAPHITE variants compared to other XAI methods, such as Attention (AUROC = 0.92) and FullGrad (AUROC = 0.91). Model-agnostic methods LIME (AUROC = 0.83) and SHAP (AUROC = 0.61) show lower performance in this context. The colour gradient along the curves represents different threshold values.
Figure 5: Precision-recall (PR) curves for various XAI methods on breast cancer TMA analysis. AUPRC for each method is displayed in the legend, with GRAPHITE-V2 achieving the highest AUPRC of 0.78, followed by GRAPHITE-Base and Attention, both with an AUPRC of 0.77. This performance highlights the superior PR balance of GRAPHITE variants compared to other XAI methods, such as FullGrad (AUPRC = 0.65) and Grad-CAM++ (AUPRC = 0.62). LIME (AUPRC = 0.52) and SHAP (AUPRC = 0.31) exhibit significantly lower precision-recall performance. The colour gradient along the curves represents different threshold values.
Figure 6: Comparison of XAI methods based on the CXPS metric. The GRAPHITE variants (GRAPHITE-Base, GRAPHITE-V1, GRAPHITE-V2) achieve the highest CXPS scores, with GRAPHITE-V2 scoring 0.65, indicating its superior balance of interpretability and predictive performance. Other methods, such as Attention and FullGrad, perform moderately well, while EigenCAM, AblationCAM and GradCAM achieve comparatively lower CXPS scores. Model-agnostic methods LIME (CXPS=0.48) and particularly SHAP (CXPS=0.35) score lowest, reflecting limitations in overall performance and interpretability coherence in this application according to our composite metric. This evaluation highlights GRAPHITE’s advantage in delivering a reliable and interpretable solution for breast cancer TMA analysis.
@@ 3.7 Comparison of GRAPHITE Variants
Figure 8: Part 1 of 2. Comparative visualization of breast cancer TMA cores using GRAPHITE, attention-based methods, gradient-based methods (GradCAM, GradCAM++, FullGrad), and model-agnostic methods (LIME and SHAP), alongside the original TMA core images and pathologist-annotated masks. Each row corresponds to a different TMA core. GRAPHITE demonstrates improved alignment with the annotated masks, enhancing interpretability by better capturing relevant tumour regions. In contrast, SHAP visualizations appear more diffuse and less focused on primary tumour areas.
@@ 3.8 Comparison With Traditional XAI Methods
Gradient-based methods, such as GradCAM and its advanced variant, Grad-CAM++, which utilize class-specific gradient information, achieved moderate AUROCs of 0.86 and 0.89 respectively, demonstrating their foundational class-discriminative capabilities. FullGrad, which extends gradient-based approaches by incorporating both input and neuron sensitivity components, showed improved performance (AUROC: 0.91) but exhibited limitations in localization precision metrics (mAP: 0.52, mIoU: 0.34). EigenCAM’s principal component analysis approach and AblationCAM’s feature removal strategy demonstrated comparatively lower performance across all metrics, suggesting potential limitations in their underlying theoretical frameworks for this complex task.
Attention-based methods, like MIL-Attention (representing the attention mechanism within the classifier itself), establish a competitive baseline (AUROC: 0.92, AUPRC: 0.77, ThR: 0.60), matching GRAPHITE-Base’s performance. Model-agnostic methods LIME and SHAP were also evaluated. LIME achieved moderate results (AUROC: 0.83, AUPRC: 0.52), outperforming basic GradCAM but falling short of more advanced methods. SHAP, in contrast, demonstrated limited performance across most metrics (AUROC: 0.61, AUPRC: 0.31, mIoU: 0.15). This suboptimal performance may be attributed to inherent limitations in SHAP’s DeepExplainer, which is built upon the DeepLIFT algorithm. Although, DeepLIFT introduces rules (such as the RevealCancel rule and difference-based propagation) to manage non-linearities and saturation, its assumptions may not hold in the presence of complex network behaviours[78]. This may lead to oversimplified explanations that fail to capture spatial dependencies and non-linear patterns intrinsic to histopathological images. These results suggest that while model-agnostic methods offer flexibility, they may not capture the nuanced, spatially correlated features critical for histopathology interpretation.
GRAPHITE-V1 advances the discriminative capability over the base attention with an AUROC of 0.93, though with a marginally lower AUPRC (0.76). GRAPHITE-V2 ultimately achieves superior performance across all metrics (AUROC: 0.94, AUPRC: 0.78, ThS: 0.50, ThR: 0.70), highlighting the effectiveness of its integrated architectural innovations (multiscale graph structure, GAT, SAN, and confidence-based fusion) in delivering both reliable predictions and interpretable visualizations (Figures 8 and 9).
Figure 9: Part 2 of 2. Continuation of comparative visualizations from Figure 8, showing additional TMA core examples analyzed using the same set of interpretation techniques. Observations are consistent with Figure 8: GRAPHITE continues to show better alignment with pathologist annotations, while SHAP remains less localized in identifying tumour regions.
Figure 10: Net Benefit Curves for various XAI methods, illustrating the clinical utility of each method across a range of decision thresholds. The net benefit is plotted as a function of threshold, showing the trade-off between true positives and the cost of false positives for each method. The AUDC is provided in the legend for each method as a quantitative measure of the model’s utility. GRAPHITE-V2 achieves the highest AUDC (4.16e+5), followed by GRAPHITE-V1 (4.08e+5) and GRAPHITE-Base (4.15e+5), indicating superior net benefit across thresholds. These results highlight the advantage of the GRAPHITE variants over other XAI methods, such as Attention (AUDC = 4.04e+5) and FullGrad (AUDC = 3.55e+5).SHAP (AUDC = 1.72e+5) show considerably lower clinical utility according to this analysis. The dashed line represents the “no net benefit” baseline, where a model yields no additional benefit over treating all or none of the cases. The higher net benefit values for GRAPHITE variants underscore their potential effectiveness in clinical decision-making for breast cancer analysis.